OpenStudy (anonymous):

If a = b + c and c ___ 0, then a > b. Choose the relationship symbol that makes a true statement. < = >

OpenStudy (anonymous):

@elementwielder @Loser66

geerky42 (geerky42):

What do you think?

OpenStudy (anonymous):

=?

geerky42 (geerky42):

hmm, if c = 0, then a = b + c would be a = b + 0 -> a = b

geerky42 (geerky42):

not what we want. we need a > b

geerky42 (geerky42):

so we eliminated = Which is it? > or < ?

OpenStudy (anonymous):

use an example. say b is 1 and c is 1, meaning c is greater than (>) 0. that would mean a would be 2, which is greater than b. if we were to say b is 1 and c is -1, meaning c is less than 0. a would then have to be 0, which is less than b. does that make sense?

OpenStudy (anonymous):

ok thnx